# Bridge multisections of knotted surfaces in $S^{4}$

### v1

@@ Abstract.
Bridge multisections are combinatorial descriptions of surface links in 4-space using tuples of trivial tangles. They were introduced by Islambouli, Karimi, Lambert-Cole, and Meier to study curves in rational surfaces. In this paper, we prove a uniqueness result for bridge multisections of surfaces in 4–space: we give a complete set of moves relating to any two multiplane diagrams of the same surface. This is done by developing a surgery operation on multiplane diagrams called band surgery. Another application of this surgery move is that any $n$–valent graph with an $n$–edge coloring is the spine of a bridge multisection for an unknotted surface. We also prove that any multisected surface in $S^{4}$ can be unknotted by finitely many band surgeries.
@@ 1. Introduction
Embedded surfaces in 4-space can be described using movies, marked vertex diagrams, broken surface diagrams, or via braid charts [CKS04a]. Some of these presentations resemble techniques for studying knots in 3-space. For instance, the 4-dimensional analog of a knot projection is a broken surface diagram, where the self-intersections of an immersed surface in 3-space are decorated with crossing information.
Bridge splittings for links in 3–space, often called plat projections, are classic tools in knot theory [AFH+21, Ch 27]. In 2017, Meier and Zupan introduced an analog of plats for surface-links in 4–space [MZ17]. Triplane diagrams are triplets of trivial tangles such that the union of any two is a 3–dimensional bridge presentation of an unlink in $S^{3}$. Any surface can be described using a triplane diagram. Since their introduction, triplane diagrams have been used to study known surface-link invariants such as colorings [ST22], Seifert solids [JMMZ22b], triple point numbers [JMMZ22a], and the fundamental group of the complement [JP23]. Examples of new invariants obtained from triplane diagrams are the crossing number [AAD+23], group trisections [BKK+24], Nielsen equivalence [JMMZ22a], and $\mathcal{L}$-invariants [BCTT22, APZ23].
In 2022, Islambouli, Karimi, Lambert-Cole, and Meier introduced a generalization of triplane diagrams where the number of tangles is arbitrary [IKLCM22]. A multiplane diagram $\mathcal{T}$ is a tuple $(T_{1},\dots,T_{n})$ of trivial tangles such that consecutive pairs, including the last and first tangles, form bridge presentations of unlinks in 3–space. As with triplanes, multiplane diagrams determine embeddings of surfaces in the 4–sphere. Invariants of triplane diagrams, such as the $\mathcal{L}$-invariant [ABG+23], have been generalized to this setting. In Section 3, we survey the known families of multiplane diagrams in the literature, such as 2–bridge multisections, completely decomposable diagrams, and multiplane diagrams of twist spun knots.
One advantage of multiplane diagrams is that some complexities of the multisected surface, like the bridge number, can be reduced at the expense of letting the number of tangles in a diagram be large. For instance, in Section 3.1, we show that every unknotted surface admits a multisection with 2–bridge tangles, while the Euler characteristic of surfaces which can be represented by $b$–bridge triplane diagrams for a given $b$ is bounded.
The main contribution of this paper is a calculus for multiplane diagrams of embedded surfaces in 4–space. Meier and Zupan introduced a complete set of triplane moves, relating any two triplane diagrams representing isotopic surfaces in 4–space. In Section 5, we expand the triplane moves from [MZ17] to a sufficient set of multiplane moves.
@@
Let $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ be two multiplane diagrams representing isotopic surfaces in 4–space. There is a finite sequence of multiplane moves that turns $\mathcal{T}_{1}$ into $\mathcal{T}_{2}$.
@@
Figure 1. Turning a $(1;1)$–bridge 4–section of an unknotted 2–sphere into a $(2;1)$–bridge 4–section of a Klein bottle.
Figure 2. Turning a $(2;2,2,2,2)$–bridge quadrisection of a 2–component unlink of 2–spheres into a $(3;2,2,2)$–bridge trisection of a spun trefoil.
In Section 6, we prove any 1–handle addition on a multisected surface can be achieved by a sequence of band surgeries. Thus, band surgery is an unknotting move for multiplane diagrams, just as 1–handle addition can unknot any surface in 4–space [HK79].
@@
As a consequence of Theorems 1.1 and 1.2, we obtain that any two multiplane diagrams are related by multiplane moves, band surgeries, and their inverses (see Theorem 6.3). This motivates the problem of using band surgeries on multiplane diagrams to find bounds for the unknotting numbers of 2–knots discussed in [JKRS21].
Another application of band surgery is related to the graph isomorphism class of the spine of a bridge multisection. The spine of a multiplane diagram is the graph obtained by taking the endpoint union of its tangles. This graph is regular and equipped with edge coloring induced by the tangles in the given multiplane. In [MTZ23], Meier, Thompson, and Zupan proved that every 3–valent graphs with 3–edge colorings arise as the spine of a bridge trisection for unknotted surfaces. In Section 7, we prove the analogous result for $n$–valent graphs.
@@
Let $\Gamma$ be an $n$–valent graph with an $n$–edge coloring. There is a bridge multisection $\mathcal{T}$ for an unknotted surface in $S^{4}$ with 1-skeleton isomorphic to $\Gamma$ and inducing the edge coloring of $\Gamma$.
One way to interpret Theorem 1.3 is that the isomorphism type of the spine of a multiplane diagram does not detect the knottedness of the underlying surface. On the other hand, by keeping track of the crossing information of each tangle of the spine, one can build distinct multiplane diagrams for the same knotted surface. In Section 3.5, we show that any knotted surface in 4–space has infinitely many distinct multisections with at least four sectors.
@@
Suppose that $F\subset S^{4}$ admits a $b$–bridge $n$–section for some $b\geq 2$, $n\geq 4$. Then for any $b^{\prime}\geq b+2$, $F$ admits infinitely many non-diffeomorphic $b^{\prime}$–bridge $n$–sections.
A consequence of Theorem 1.4 is that the unknotted 2–sphere admits non-standard bridge multisections with at least four sectors, and so Otal’s theorem does not hold for bridge multisections [Ota82]. More precisely, not every multisection of the unknotted 2–sphere is perturbed. The existence of non-standard multisections was first observe in [IKLCM22, Thm 2.5] where the authors found infinitely many non-diffeomorphic 3–bridge 4–sections of the unknotted sphere. It is important to note that the existence of non-standard bridge trisections of the unknotted 2–sphere is still unknown.
@@ 2. Background
We are interested on smooth embeddings of closed surfaces into $S^{4}$. Following [HK79], an embedded surface is said to be unknotted if it can be decomposed as the distant sum of connected sums of orientable unknotted surfaces and standard projective planes. Unknotted orientable surfaces are those that bound smoothly embedded 3–dimensional handlebodies. Equivalently, an orientable surface is unknotted if it can be isotoped to be embedded into $S^{3}\subset S^{4}$. There are two types of standard projective planes, $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$, both differentiated by their Euler number. Figure 3 depicts movies of $\mathcal{P}_{\pm}$.
@@
Given an embedded surface $F\subset S^{4}$, the crosscap sum of $F$ is the new surface obtained by adding a $\mathcal{P}_{\pm}$ summand to $F$. We say that $F^{\prime}$ is obtained from $F$ by 1–handle addition if $F^{\prime}$ is the result of replacing two small disks in $F$ with a thin tube connecting them [CKS04b]. More precisely, if $D^{2}\times I$ is an embedded 1–handle for $F$, so that $F\cap(D^{2}\times I)=D^{2}\times\{0,1\}$, 1–handle addition of $F$ is the new surface $F^{\prime}=F\setminus\left(D^{2}\times\{0,1\}\right)\cup\left(S^{1}\times I\right)$. We refer to the core of the 1–handle $\rho=\{0\}\times I$ as the guiding arc of the 1–handle. The property of being unknotted is preserved by 1–handle addition.
@@ 2.1. Trivial tangles
We will work with boundary parallel $(n-2)$–balls properly embedded in $B^{n}$. In dimension four, a collection of properly embedded disks $\mathcal{D}\subset B^{4}$ is called trivial if $\mathcal{D}$ is isotopic into the boundary of the 4–ball while fixing $\partial\mathcal{D}$. If $|\mathcal{D}|=c$, then we may refer to $\mathcal{D}$ as a $c$–patch trivial disk system. The following result implies that $c$–trivial disk systems are unique up to isotopy.
@@
In dimension three, a trivial tangle is a tuple $(B,T)$, or simply $T\subset B$, where $B$ is a 3-ball containing properly embedded arcs $T$ such that, fixing the endpoints of $T$, we may isotope $T$ into $\partial B$. We consider the endpoints of a tangle to be punctures of the 2–sphere $\Sigma=\partial B$. Given a trivial tangle $T\subset B$, one can find $|T|$ embedded disks in $B$ with boundary equal to the union of $T$ with arcs in $\Sigma$. Such a collection of disks is called a set of bridge disks for $T$. For a link $L\subset S^{3}$, a decomposition $(S^{3},L)=(B_{1},T_{1})\cup_{\Sigma}(B_{2},T_{2})$, where each pair $(B_{i},T_{i})$ is a trivial tangle, is called a bridge splitting. If each tangle $T_{i}$ has $b$ strands, we sometimes call this decomposition a $b$–bridge splitting of $L$. The symbol $\overline{T}$ denotes the mirror image of a trivial tangle $T$.
@@ 2.2. Bands
Fix $T\subset B$ a properly embedded 1–manifold in a 3-ball. Classically, a band is an embedding of a rectangle that intersects a link on two opposite sides [Miy86]. In this work, we will use a broader definition of bands (and band surgery) to consider rectangles that interact with the boundary of $B$ (see Figure 4). More precisely, a band $v$ is an embedding of a rectangle $v:I\times I\rightarrow B$ with $(T\cup\Sigma)\cap Im(v)\subset v(I\times\{0,1\})$ and satisfying one of the following conditions:
@@
In each case, we can define a new tangle $T[v]$ as in Figure 4. Note that in cases (b) or (c), surgery along $v$ turns a $b$–stranded trivial tangle $T$ into a tangle with $(b+1)$ strands. To define band surgery in Section 4.2, we will consider bands as in (b) and (c) where the cores are monotonic arcs (no critical point). Under these assumptions, $T[v]$ will be a trivial tangle.
Figure 4. Three kinds of bands.
@@
In Section 4.2, it will be useful to describe a band as in case (b) or (c) using a pair $(\alpha,f)$ where $\alpha$ is an embedded arc with framing $f$. Suppose $T$ and $T^{\prime}$ are trivial tangles with the same endpoints. If $(\alpha,f)$ represents a band for a properly embedded 1–manifold $L$, we denote by $L[\alpha,f]$ the result of $L$ after banded surgery. If $(\alpha,f)$ and $(\alpha^{\prime},f^{\prime})$ are framed arcs in $B$ and $B^{\prime}$, respectively, with the same endpoint $*\in int(\Sigma)$, we denote by $f+f^{\prime}$ the new framing on $\alpha\cup\alpha^{\prime}$ provided the $f$ and $f^{\prime}$ agree at *. This way, $(\alpha\cup\alpha^{\prime},f+f^{\prime})$ will represent a band for the link $T\cup\overline{T}^{\prime}$ intersecting the bridge sphere in one arc. See Figure 8 for an example of this situation.
@@
Intuitively, after a diffeomorphism of the 3–ball, the bridge disks and bands look like in Figure 5. Dual bands have the property that the surgered $b$–stranded tangle $T[v]$ is also trivial. An equivalent definition of dual bands using shadow arcs instead of bridge disks can be found in [MZ17, Sec 3]. The following result is contained in the proof of Lemma 3.3 of [MZ17].
@@
Let $T_{1}$ and $T_{2}$ be two $b$–bridge trivial tangles with the same endpoints. If $T_{1}\cup\overline{T}_{2}$ is a $c$–component unlink, then there exist a set $v$ of $(b-c)$ bands such that $v$ is dual to $T_{1}$, and $T_{1}[v]=T_{2}$.
@@ 3. Bridge multisections
We discuss a 4-dimensional analog of bridge position for links in $S^{3}$. Bridge multisections can be defined for surfaces in arbitrary 4-manifolds (see [IKLCM22] or [ABG+23]). We write a simplified version for $F\subset S^{4}$ as this paper only studies surfaces in 4–space.
Definition 3.1 ([IKLCM22]).
Let $F$ be an embedded surface in $S^{4}$. A $(b;c)$–bridge multisection of $F$, where $c=(c_{i})_{i=1}^{n}$, is a decomposition
@@
    $(B_{i},T_{i})=(X_{i},\mathcal{D}_{i})\cap(X_{i-1},\mathcal{D}_{i-1})$ is a $b$–bridge trivial tangle inside a 3-ball, and
@@
    $(\Sigma,\{p_{1},\dots,p_{2b}\})=\bigcap_{i=1}^{n}(X_{i},\mathcal{D}_{i})$ is a 2–sphere with $2b$–punctures.
The decomposition $S^{4}=\bigcup_{i=1}^{n}X_{i}$ is called the 0–multisection of $S^{4}$. In words, a $(b;c)$–bridge multisection, also referred to as $b$–bridge $n$–section, is a decomposition of the surface $F$ into trivial disk systems intersecting each 3–ball $B_{i}=X_{i}\cap X_{i-1}$ in $b$–bridge trivial tangles $T_{i}$. The surface $\Sigma$ is called the central surface or bridge surface of $F$ and it is often depicted as a $2b$–punctured sphere. The numbers $c_{1},\dots,c_{n}$ are called the patch numbers, and $b=|\Sigma\cap F|/2$ is called the bridge number of the decomposition. When all patch numbers are equal, say to $c_{1}$, we will write $c$ as a number. For instance $(4;2)$–bridge multisections are $\left(4;(2,\dots,2)\right)$–bridge multisections. A $(b;c)$–bridge multisection is thin if $c_{i}=b-1$ for all $i\in\mathbb{Z}_{n}$. Theorem 5.6 states that every surface admits a thin multisection.
The union of the tangles $\bigcup_{i=1}^{n}(B_{i},T_{i})$ is called the spine of the bridge multisection. Abstractly, the spine is a $n$–valent graph with $2b$ vertices. It follows from Definition 3.1 that the union of consecutive tangles $L_{i}=T_{i}\cup\overline{T}_{i+1}$ is a $c_{i}$-component unlink in $S^{3}$. By Lemma 2.2, there is a unique way to fill the unlink with a trivial disk tangle in the 4–ball. Thus, after fixing the embeddings of the 3–balls $\{B_{i}\}_{i=1}^{n}$ in $S^{4}$, the tuple of tangles $(T_{1},\dots,T_{n})$ determines a unique embedding of $F$ up to isotopy.
@@
A multiplane diagram, also called $n$–plane diagram, is an ordered tuple of trivial tangles with the same endpoints $\mathcal{T}=(T_{1},\dots,T_{n})$ such that, for each $i\in\mathbb{Z}_{n}$, $T_{i}\cup\overline{T}_{i+1}$ is an unlink in $S^{3}$.
@@
Two bridge multisections $(S^{4},F)=\bigcup_{i=1}^{n}(X_{i},\mathcal{D}_{i})$ and $(S^{4},F^{\prime})=\bigcup_{i=1}^{n}(X^{\prime}_{i},\mathcal{D}^{\prime}_{i})$ are said to be diffeomorphic if there is a diffeomorphism $\varphi:(S^{4},F)\rightarrow(S^{4},F^{\prime})$ satisfying $\varphi(X_{i},\mathcal{D}_{i})=(X_{\sigma(i)},\mathcal{D}_{\sigma(i)})$ for all $i\in\mathbb{Z}_{n}$, for some permutation $\sigma$. If $\varphi$ is isotopic to the identity on $S^{4}$, we say that the bridge multisections are isotopic. One way to obstruct multiplane diagrams from giving diffeomorphic bridge multisections is by looking at the links obtained by gluing two non-adjacent tangles. This approach was exploited in [IKLCM22, Thm 2.5] to build non-standard 4-sections of the unknotted sphere. Given a multiplane diagram $\mathcal{T}$, a cross-section of $\mathcal{T}$ is a link in $S^{3}$ isotopic to $T_{i}\cup\overline{T}_{j}$ for some $i,j\in\mathbb{Z}_{n}$.
@@
Let $\mathcal{T}$ and $\mathcal{T}^{\prime}$ be two multiplane diagrams. Assume that the induced bridge multisections of $F_{\mathcal{T}}$ and $F_{\mathcal{T}^{\prime}}$ are diffeomorphic. Then, after possibly permuting the tangles in $\mathcal{T}^{\prime}$, the $ij$–cross-sections of $\mathcal{T}$ and $\mathcal{T}^{\prime}$ differ by a diffeomorphism of $S^{3}$.

### v2

@@ Abstract.
Bridge multisections are combinatorial descriptions of surface links in $S^{4}$ using tuples of trivial tangles. They were introduced by Islambouli, Karimi, Lambert-Cole, and Meier to study curves in rational surfaces. In this paper, we prove a uniqueness result for bridge multisections of surfaces in $S^{4}$: we give a complete set of moves relating to any two multiplane diagrams of the same surface. This is done by developing a surgery operation on multiplane diagrams called band surgery. Another application of this surgery move is that any nn–valent graph with an nn–edge coloring is the spine of a bridge multisection for an unknotted surface. We also prove that any multisected surface in $S^{4}$ can be unknotted by finitely many band surgeries.
@@ 1. Introduction
Smoothly embedded surfaces in $S^{4}$ can be described using movies, marked vertex diagrams, broken surface diagrams, or via braid charts [CKS04]. Some of these presentations resemble techniques for studying knots in 3–space. For instance, the 4-dimensional analog of a knot projection is a broken surface diagram, where the self-intersections of an immersed surface in 3–space are decorated with crossing information.
Bridge splittings for links in 3–space, often called plat projections, are classic tools in knot theory [AFH+02, Ch 27]. In 2017, motivated by Gay and Kirby’s theory of trisections of 4–manifolds [GK16], Meier and Zupan introduced an analog of plats for surface-links in $S^{4}$ [MZ17]. Triplane diagrams are triplets of trivial tangles such that the union of any two is a 3–dimensional bridge presentation of an unlink in $S^{3}$. Any surface can be described using a triplane diagram. Since their introduction, triplane diagrams have been used to study known surface-link invariants such as colorings [ST22], Seifert solids [JMM+22b], triple point numbers [JMM+22a], and the fundamental group of the complement [JP23]. Examples of new invariants obtained from triplane diagrams are the crossing number [AAD+23], group trisections [BKK+24], Nielsen equivalence [JMM+22a], and $\mathcal{L}$-invariants [BCT+22, APZ23b].
In 2022, Islambouli, Karimi, Lambert-Cole, and Meier introduced a generalization of triplane diagrams where the number of tangles is arbitrary [IKL+22]. A multiplane diagram $\mathcal{T}$ is a tuple $(T_{1},\dots,T_{n})$ of trivial tangles such that consecutive pairs, including the last and first tangles, form bridge presentations of unlinks in 3–space. As with triplanes, multiplane diagrams determine embeddings of surfaces in the 4–sphere. Invariants of triplane diagrams, such as the $\mathcal{L}$-invariant [ABG+23a], have been generalized to this setting. In Section 3, we survey the known families of multiplane diagrams in the literature, such as 2–bridge multisections, completely decomposable diagrams, and multiplane diagrams of twist spun knots.
One advantage of multiplane diagrams is that some complexities of the multisected surface, like the bridge number, can be reduced at the expense of letting the number of tangles in a diagram be large. For instance, in Section 3.1, we show that every unknotted surface admits a multisection with 2–bridge tangles, while the Euler characteristic of surfaces which can be represented by bb–bridge triplane diagrams for a given $b$ is bounded.
The main contribution of this paper is a calculus for multiplane diagrams of smoothly embedded surfaces in $S^{4}$. Meier and Zupan introduced a complete set of triplane moves, relating any two triplane diagrams representing isotopic surfaces in $S^{4}$. In Section 5, we expand the triplane moves from [MZ17] to a sufficient set of multiplane moves.
@@
Let $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ be two multiplane diagrams representing isotopic surfaces in $S^{4}$. There is a finite sequence of multiplane moves that turns $\mathcal{T}_{1}$ into $\mathcal{T}_{2}$.
@@
Figure 1. Turning a (1;1)(1;1)–bridge 4–section of an unknotted 2–sphere into a (2;1)(2;1)–bridge 4–section of a Klein bottle.
Figure 2. Turning a (2;2,2,2,2)(2;2,2,2,2)–bridge quadrisection of a 2–component unlink of 2–spheres into a (3;2,2,2,2)(3;2,2,2,2)–bridge quadrisection of a spun trefoil.
In Section 6, we prove any 1–handle addition on a multisected surface can be achieved by a sequence of band surgeries. Thus, band surgery is an unknotting move for multiplane diagrams, just as 1–handle addition can unknot any smooth surface in $S^{4}$ [HK79].
@@
As a consequence of Theorems 5.10 and 6.2, we obtain that any two multiplane diagrams are related by multiplane moves, band surgeries, and their inverses (see Theorem 6.3). This motivates the problem of using band surgeries on multiplane diagrams to find bounds for the unknotting numbers of 2–knots discussed in [JKR+21]. Question 1.1 is motivated by the fact that the bridge index of knots in $S^{3}$ gives an upper bound to their tunnel number.
Question 1.1.
Is there a relationship between the stabilization number of a surface $F\subset S^{4}$ and the bridge/patch numbers of a bridge multisection of $F$?
Another application of band surgery is related to the graph isomorphism class of the spine of a bridge multisection. The spine of a multiplane diagram is the graph obtained by taking the endpoint union of its tangles. This graph is regular and equipped with edge coloring induced by the tangles in the given multiplane. In [MTZ23], Meier, Thompson, and Zupan proved that every 3–valent graph with a 3–edge coloring arises as the spine of a bridge trisection for an unknotted surface. In Section 7, we prove the analogous result for nn–valent graphs.
@@
Let $\Gamma$ be an nn–valent graph with an nn–edge coloring. There is a bridge multisection $\mathcal{T}$ for an unknotted surface in $S^{4}$ with 1–skeleton isomorphic to $\Gamma$ and inducing the edge coloring of $\Gamma$.
One way to interpret Theorem 7.2 is that the isomorphism type of the spine of a multiplane diagram does not detect the knottedness of the underlying surface. On the other hand, by keeping track of the crossing information of each tangle of the spine, one can build distinct multiplane diagrams for the same knotted surface. In Section 3.5, we show that any knotted surface in $S^{4}$ has infinitely many distinct multisections with at least four sectors.
@@
Suppose that $F\subset S^{4}$ admits a bb–bridge nn–section for some $b\geq 2$, $n\geq 4$. Then for any $b^{\prime}\geq b+2$, $F$ admits infinitely many non-diffeomorphic $b^{\prime}$–bridge nn–sections.
A consequence of Theorem 3.13 is that the unknotted 2–sphere admits non-standard bridge multisections with at least four sectors, and so Otal’s theorem does not hold for bridge multisections [OTA82]. More precisely, not every multisection of the unknotted 2–sphere is perturbed. The existence of non-standard multisections was first observed in [IKL+22, Thm 2.5] where the authors found infinitely many non-diffeomorphic 3–bridge 4–sections of the unknotted sphere. It is important to note that the existence of non-standard bridge trisections of the unknotted 2–sphere is still unknown.
@@ 2. Background
We are interested in smooth embeddings of closed surfaces into $S^{4}$; even when not stated, smoothness will be assumed throughout this work. Following [HK79], an embedded surface is said to be unknotted if it can be decomposed as the distant sum [^1] See Section 3.3 for the defition. of connected sums of orientable unknotted surfaces and standard projective planes. Unknotted orientable surfaces are those that bound smoothly embedded 3–dimensional handlebodies. Equivalently, an orientable surface is unknotted if it can be isotoped to be embedded into $S^{3}\subset S^{4}$. There are two types of standard projective planes, $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$, both differentiated by their Euler number; $e(\mathcal{P}_{\pm})=\pm 2$ [JMM+22a, Rem 2.6]. Figure 3 depicts movies of $\mathcal{P}_{\pm}$.
@@
Given an embedded surface $F\subset S^{4}$, the crosscap sum of $F$ is the new surface obtained by adding a $\mathcal{P}_{\pm}$ summand to $F$. We say that $F^{\prime}$ is obtained from $F$ by 1–handle addition if $F^{\prime}$ is the result of replacing two small disks in $F$ with a thin tube connecting them [CKS04]. More precisely, if $D^{2}\times I$ is an embedded 1–handle for $F$, so that $F\cap(D^{2}\times I)=D^{2}\times\{0,1\}$, 1–handle addition of $F$ is the new surface $F^{\prime}=F\setminus\left(D^{2}\times\{0,1\}\right)\cup\left(S^{1}\times I\right)$. We refer to the core of the 1–handle $\rho=\{0\}\times I$ as the guiding arc of the 1–handle. The property of being unknotted is preserved by 1–handle addition.
@@ 2.1. Trivial tangles
We will work with boundary parallel (n2)(n-2)–balls properly embedded in $B^{n}$. In dimension four, a collection of properly embedded disks $\mathcal{D}\subset B^{4}$ is called trivial if $\mathcal{D}$ is isotopic into the boundary of the 4–ball while fixing $\partial\mathcal{D}$. If $|\mathcal{D}|=c$, then we may refer to $\mathcal{D}$ as a cc–patch trivial disk system. The following result implies that cc–trivial disk systems are unique up to isotopy.
@@
In dimension three, a trivial tangle is a tuple $(B,T)$, or simply $T\subset B$, where $B$ is a 3–ball containing properly embedded arcs $T$ such that, fixing the endpoints of $T$, we may isotope $T$ into $\partial B$. We consider the endpoints of a tangle to be punctures of the 2–sphere $\Sigma=\partial B$. Given a trivial tangle $T\subset B$, one can find $|T|$ embedded disks in $B$ with boundary equal to the union of $T$ with arcs in $\Sigma$. Such a collection of disks is called a set of bridge disks for $T$. For a link $L\subset S^{3}$, a decomposition $(S^{3},L)=(B_{1},T_{1})\cup_{\Sigma}(B_{2},T_{2})$, where each pair $(B_{i},T_{i})$ is a trivial tangle, is called a bridge splitting. If each tangle $T_{i}$ has $b$ strands, we sometimes call this decomposition a bb–bridge splitting of $L$. The symbol $\overline{T}$ denotes the mirror image of a trivial tangle $T$.
@@ 2.2. Bands
Fix $T\subset B$ a properly embedded 1–manifold in a 3–ball. Classically, a band is an embedding of a rectangle that intersects a link on two opposite sides [MIY86]. In this work, we will use a broader definition of bands (and band surgery) to consider rectangles that interact with the boundary of $B$ (see Figure 4).
Definition 2.3.
A band $v$ is an embedding of a rectangle $v:I\times I\rightarrow B$ with $(T\cup\Sigma)\cap Im(v)\subset v(I\times\{0,1\})$ and satisfying one of the following conditions:
@@
In each case, we define the band surgery of $T$ along $v$, denoted by $T[v]$, to be a new tangle as in Figure 4.
Note that in cases (b) or (c), surgery along $v$ turns a bb–stranded trivial tangle $T$ into a tangle with $(b+1)$ strands. To define band surgery in Section 4.2, we will consider bands as in (b) and (c) where the cores are monotonic arcs (no critical point). Under these assumptions, $T[v]$ will be a trivial tangle.
Figure 4. (Top) Three kinds of bands $v$ with sides on a tangle $T$: (a) classical, (b) type 1, and (c) type 0. (Bottom) The band surgered tangle $T[v]$.
@@
In Section 4.2, it will be useful to describe a band as in case (b) or (c) using a pair $(\alpha,f)$ where $\alpha$ is an embedded arc with framing $f$. In such cases, we will represent bands as framed arrows, where the arrow indicates the direction in which the tangle will be moved; see Figure 1. When the framing is not explicit, we will consider the blackboard framing as in Figure 2.
Suppose $T$ and $T^{\prime}$ are trivial tangles with the same endpoints. If $(\alpha,f)$ represents a band for a properly embedded 1–manifold $L$, we denote by $L[\alpha,f]$ the result of $L$ after banded surgery. If $(\alpha,f)$ and $(\alpha^{\prime},f^{\prime})$ are framed arcs in $B$ and $B^{\prime}$, respectively, with the same endpoint $\in int(\Sigma)$, we denote by $f+f^{\prime}$ the new framing on $\alpha\cup\alpha^{\prime}$ provided the $f$ and $f^{\prime}$ agree at *. This way, $(\alpha\cup\alpha^{\prime},f+f^{\prime})$ will represent a band for the link $T\cup\overline{T}^{\prime}$ intersecting the bridge sphere in one arc. See Figure 8 for an example of this situation.
@@
Intuitively, after a diffeomorphism of the 3–ball, the bridge disks and bands look like Figure 5. An equivalent definition of dual bands using shadow arcs instead of bridge disks can be found in [MZ17, Sec 3]. Dual bands have the property that the surgered bb–stranded tangle $T[v]$ is also trivial; see [MZ17, Lem 3.1]. The following result is contained in the proof of Lemma 3.3 of [MZ17].
@@
Let $(B_{1},T_{1})$ and $(B_{2},T_{2})$ be two bb–bridge trivial tangles with the same endpoints. If $T_{1}\cup\overline{T}_{2}$ is a cc–component unlink in $S^{3}=B_{1}\cup\overline{B}_{2}$, then there exist a set $v$ of (bc)(b-c) bands such that $v$ is dual to $T_{1}$, and $T_{1}[v]=T_{2}$.
@@ 3. Bridge multisections
We discuss a 4-dimensional analog of bridge position for links in $S^{3}$. Bridge multisections can be defined for surfaces in arbitrary 4-manifolds (see [IKL+22] or [ABG+23a]). We write a simplified version for $F\subset S^{4}$ as this paper only studies surfaces in $S^{4}$. The curious reader may find interesting the work of Islambouli and Naylor on multisections of closed 4–manifolds [IN24].
Definition 3.1 ([IKL+22]).
Let $F$ be an embedded surface in $S^{4}$. A (b;c)(b;c)–bridge multisection of $F$, where $c=(c_{i})_{i=1}^{n}$, is a decomposition
@@
    $(B_{i},T_{i})=(X_{i},\mathcal{D}_{i})\cap(X_{i-1},\mathcal{D}_{i-1})$ is a bb–bridge trivial tangle inside a 3–ball, and
@@
    $(\Sigma,\{p_{1},\dots,p_{2b}\})=\bigcap_{i=1}^{n}(X_{i},\mathcal{D}_{i})$ is a 2–sphere with 2b2b–punctures.
The decomposition $S^{4}=\bigcup_{i=1}^{n}X_{i}$ is called the 0–multisection of $S^{4}$. In words, a (b;c)(b;c)–bridge multisection, also referred to as bb–bridge nn–section, is a decomposition of the surface $F$ into trivial disk systems intersecting each 3–ball $B_{i}=X_{i}\cap X_{i-1}$ in bb–bridge trivial tangles $T_{i}$. The surface $\Sigma$ is called the central surface or bridge surface of $F$ and it is often depicted as a 2b2b–punctured sphere. The numbers $c_{1},\dots,c_{n}$ are called the patch numbers, and $b=|\Sigma\cap F|/2$ is called the bridge number of the decomposition. When all patch numbers are equal, say to $c_{1}$, we will write $c$ as a number. For instance (4;2)(4;2)–bridge multisections are $\left(4;(2,\dots,2)\right)$–bridge multisections. A (b;c)(b;c)–bridge multisection is thin if $c_{i}=b-1$ for all $i\in\mathbb{Z}_{n}$. Theorem 5.6 states that every surface admits a thin multisection.
The union of the tangles $\bigcup_{i=1}^{n}(B_{i},T_{i})$ is called the spine of the bridge multisection. Abstractly, the spine is a nn–valent graph with $2b$ vertices. It follows from Definition 3.1 that the union of consecutive tangles $L_{i}=T_{i}\cup\overline{T}_{i+1}$ is a $c_{i}$-component unlink in $S^{3}$. By Lemma 2.2, there is a unique way to fill the unlink with a trivial disk tangle in the 4–ball. Thus, after fixing the embeddings of the 3–balls $\{B_{i}\}_{i=1}^{n}$ in $S^{4}$, the tuple of tangles $(T_{1},\dots,T_{n})$ determines a unique embedding of $F$ up to isotopy.
@@
A multiplane diagram, also called nn–plane diagram, is an ordered tuple of trivial tangles with the same endpoints $\mathcal{T}=(T_{1},\dots,T_{n})$ such that, for each $i\in\mathbb{Z}_{n}$, $T_{i}\cup\overline{T}_{i+1}$ is an unlink in $S^{3}$.
@@
An nn–plane diagram with $n=3$ is called a triplane diagram [MZ17]. They can be used to describe classical invariants of the knotted surface, such as the Euler number, the fundamental group of the complement, and the peripheral subgroup [JMM+22a]. One could study how these procedures extend to arbitrary multiplane diagrams. For instance, the natural extension of the formula for the Euler number of a triplane diagram also holds for nn–planes; the proof is the same as in [JMM+22a].
Lemma 3.3 (Cor 3.8 of [JMM+22a]).
Let $\mathcal{T}=(T_{1},\dots,T_{n})$ be an nn–plane diagram of a surface $F\subset S^{4}$. Let $w_{i}$ be the writhe of the diagram $T_{i}\cup\overline{T}_{i+1}$, where $i\in\mathbb{Z}_{n}$. Then the Euler number of $F$ is given by $e(F)=\sum_{i=1}^{n}w_{i}$.
Two bridge multisections $(S^{4},F)=\bigcup_{i=1}^{n}(X_{i},\mathcal{D}_{i})$ and $(S^{4},F^{\prime})=\bigcup_{i=1}^{n}(X^{\prime}_{i},\mathcal{D}^{\prime}_{i})$ are said to be diffeomorphic if there is a diffeomorphism $\varphi:(S^{4},F)\rightarrow(S^{4},F^{\prime})$ satisfying $\varphi(X_{i},\mathcal{D}_{i})=(X_{\sigma(i)},\mathcal{D}_{\sigma(i)})$ for all $i\in\mathbb{Z}_{n}$, for some permutation $\sigma$. If $\varphi$ is isotopic to the identity on $S^{4}$, we say that the bridge multisections are isotopic. One way to obstruct multiplane diagrams from giving diffeomorphic bridge multisections is by looking at the links obtained by gluing two non-adjacent tangles. This approach was exploited in [IKL+22, Thm 2.5] to build non-standard 4-sections of the unknotted sphere. Given a multiplane diagram $\mathcal{T}$, a cross-section of $\mathcal{T}$ is a link in $S^{3}$ isotopic to $T_{i}\cup\overline{T}_{j}$ for some $i,j\in\mathbb{Z}_{n}$. The proof of Lemma 3.4 is left to the reader.
@@
Let $\mathcal{T}$ and $\mathcal{T}^{\prime}$ be two multiplane diagrams. Assume that the induced bridge multisections of $F_{\mathcal{T}}$ and $F_{\mathcal{T}^{\prime}}$ are diffeomorphic. Then, after possibly permuting the tangles in $\mathcal{T}^{\prime}$, the ijij–cross-sections of $\mathcal{T}$ and $\mathcal{T}^{\prime}$ differ by a diffeomorphism of $S^{3}$.